# Weight-Entanglement Meets Gradient-Based Neural Architecture Search

### v1

@@ E.1 Toy Search Spaces
E.2 Language Model
We use the AdamW optimizer in all experiments pertaining to language modelling. Other hyperparameter choices are as in Table 22.
@@ E.2 Language Model
Table 22: Configurations used in DrNAS on the Language Model Spaces
E.3 AutoFormer and OFA
We use a 50%-50% train and validation split for the CIFAR10 and CIFAR100 dataset for the cell-based spaces and a 80%-20% for the weight-entangled spaces We use the official source code of AutoFormer available at code for all the AutoFormer experiments on CIFAR10 and CIFAR100. We closely follow the AutoFormer training pipeline and search space design. AutoFormer searched on three transformer sizes Autoformer-T (tiny), AutoFormer-S (small) and Autoformer-B (base). We currently restrict ourselves to Autoformer-T. For baselines like OFA and AutoFormer we follow their respective recipes to obtain the train-valid split for ImageNet-1k. Our models were trained on 2xA100s with the same effective batch-size as AutoFormer. For MobileNetV3 from once-for-all we use the same training hyperparameters as the baseline as here (in addition to architectural parameters same as Table 23) .
E.3.1 AutoFormer Fine-tuning
C10 and C100 pretrained supernet
We fine-tune the C10 and C100 selected networks (after inheriting them from the supernet) for 500 epochs. We set the learning rate to 1e-3, the warmup epochs to 5 , the warmup learning rate to 1e-6 and the min learning rate to 1e-5. All other hyperparameters are set same as Section E.3.
@@ E.3.1 AutoFormer Fine-tuning
We follow the DeiT [Touvron et al., 2021] finetuning pipeline as used in AutoFormer. Mainly we set the epochs to 1000, the warmup-epochs to 5 , the scheduler to cosine, the mixup to 0.8, the smoothing to 0.1, the weight-decay to 1e-4, the batch-size 64, the optimizer to sgd, the learning rate to 0.01 and the warmup-lr to 0.0001 for all datasets.
E.4 NB201 and DARTS
For single-stage optimizers, the supernet was trained with four different seeds. The supernet with the best validation performance from these four was discretized to obtain the final model, which is then trained from scratch four times to obtain the results shown in the table. For two-stage methods, yet again, we train the supernet four times, and perform Random Search (RS) and Evolutionary Search (ES) on each of them. The best model obtained across all four supernets for both methods are then trained from scratch with four seeds to compute the final results. For each DrNAS we follow the same recipe as suggested by the authors across all search spaces. To accommodate for multiple training recipes, we write a configurable training pipeline. The configurations for DrNAS, and SPOS are shown in Table 23 and Table 24, respectively.
@@ E.4 NB201 and DARTS
Appendix F Optimal architectures derived
F.1 DARTS
@@ F.3.1 CIFAR10
mlp_ratio:$[4,4,4,4,4,4,4,4,4,3.5,4,4,3.5,4]$, num_heads:[4,4,4,4,4,4,4,4,4,4,4,4,4,4]44444444444444[4,4,4,4,4,4,4,4,4,4,4,4,4,4][ 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 ], num_layers: 14 embed_dim: 216
@@
mlp_ratio:$[4,4,4,4,4,4,4,4,3.5,4,4,4,3.5,3.5]$, num_heads:[4,4,4,4,4,4,4,4,4,4,4,4,4,4]44444444444444[4,4,4,4,4,4,4,4,4,4,4,4,4,4][ 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 ], depth: 14, embed_dim: 240
F.3.2 CIFAR100
@@ F.4 MobileNetV3
F.5 Toy Spaces
F.5.1 Toy cell (our best architecture)
@@
F.6 Toy conv-macro (our best architecture)
@@
Appendix G Architecture Representation Analysis
@@
Centered Kernel Alignment (CKA) [Kornblith et al., 2019] is a metric, based on the Hilbert- Schmidt Independence Criterion (HSIC). This metric is primarily designed to model similarity between representations in neural networks. In this section we study the CKA between structurally identical layers in the inherited, fine-tuned and retrained networks in the AutoFormer-T space. Precisely, the goal here is to study how similar the inherited and fine-tuned representations are to the actual model trained from scratch. Table 25 Presents the average CKA values for a fixed subsample of the CIFAR10 and CIFAR100 dataset. We observe that the representations learned by the single-stage supernet are more similar to the architecture finetuned or trained from scratch. This further emphasizes the potential issues with using the inherited weights from the supernet as proxy for search for two-stage methods as highlighted in [Xu et al., 2022].
@@
CIFAR10 CIFAR100 CIFAR10 CIFAR100
TangleNAS 0.4630 0.5853 0.57125 0.65275
SPOS+ES 0.45812 0.57932 0.569374 0.6309
SPOS+RS 0.44124 0.583 0.5797 0.638948
Table 25: CKA correlation between layers
Architecture Search-Type Loss Perplexity Params Inference Time
GPT-2 Manual 3.0772 21.69 123.59M 113.3s
TangleNAS Automated 2.9038 18.243 116.51M 102.5s
  1. 1.
    Best Practices for Releasing Code
    For all experiments you report:
    1. (a)
      Did you release code for the training pipeline used to evaluate the final architectures? Yes
    2. (b)
      Did you release code for the search space? Yes
    3. (c)
      Did you release the hyperparameters used for the final evaluation pipeline, as well as random seeds? Yes
    4. (d)
      Did you release code for your NAS method? Yes
    5. (e)
      Did you release hyperparameters for your NAS method, as well as random seeds? Yes
  2. 2.
    Best practices for comparing NAS methods
    1. (a)
      For all NAS methods you compare, did you use exactly the same NAS benchmark, including the same dataset (with the same training-test split), search space and code for training the architectures and hyperparameters for that code? Yes
    2. (b)
      Did you control for confounding factors (different hardware, versions of DL libraries, different runtimes for the different methods)? Yes
    3. (c)
      Did you run ablation studies? Yes
    4. (d)
      Did you use the same evaluation protocol for the methods being compared? Yes
    5. (e)
      Did you compare performance over time? Yes
    6. (f)
      Did you compare to random search? Yes. Black box random search is very expensive in our case, hence we use the supernet from single-path-one-shot(SPOS) for random search.
    7. (g)
      Did you perform multiple runs of your experiments and report seeds? Yes
    8. (h)
      Did you use tabular or surrogate benchmarks for in-depth evaluations? Yes
  3. 3.
    Best practices for reporting important details
    1. (a)
      Did you report how you tuned hyperparameters, and what time and resources this required? N/A. We did not tune any hyperparameters.
    2. (b)
      Did you report the time for the entire end-to-end NAS method (rather than, e.g., only for the search phase)? Yes
    3. (c)
      Did you report all the details of your experimental setup? Yes
Eer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@ F.1 Toy Search Spaces
F.2 Language Model
We use the AdamW optimizer in all experiments related to language modeling. Other hyperparameter choices are as specified in Table 23. Below are the hyperparameter settings for TangleNAS. We run experiments on TinyStories on 8 RTX-2080 GPUs and experiments on OpenWebText on 8 A6000 GPUs.
@@ F.2 Language Model
Table 23: Configurations used in DrNAS on the Language Model Spaces.
F.3 AutoFormer and OFA
We use a 50%-50% train and validation split for the CIFAR-10 and CIFAR-100 datasets for the cell-based spaces and a 80%-20% for the weight-entangled spaces. We use the official source code of AutoFormer available at code autoformer for all the AutoFormer experiments on CIFAR-10 and CIFAR-100, closely following the AutoFormer training pipeline and search space design. AutoFormer explored three transformer sizes: Autoformer-T (tiny), AutoFormer-S (small), and Autoformer-B (base). We restrict ourselves to Autoformer-T.
For baselines like OFA and AutoFormer, we follow their respective recipes to obtain the train-validation split for ImageNet-1k. Our models were trained on 2xA100s with the same effective batch-size as AutoFormer. For MobileNetV3 from Once-for-All, we use the same training hyperparameters as the baseline found here (in addition to architectural parameters same as Table 24). We run experiments on CIFAR-10, CIFAR-100, and ImageNet1-k on 4 A100 GPUs.
F.3.1 AutoFormer Fine-tuning
CIFAR-10 and CIFAR-100 pretrained supernet
We fine-tune the CIFAR-10 and CIFAR-100 selected networks (after inheriting them from the supernet) for 500 epochs. We set the learning rate to 1e-3, the warmup epochs to 5, the warmup learning rate to 1e-6, and the minimum learning rate to 1e-5. All other hyperparameters are set the same as in Appendix F.3.
@@ F.3.1 AutoFormer Fine-tuning
We follow the DeiT (touvron-icml21) fine-tuning pipeline as used in AutoFormer, to finetune on downstream tasks. Specifically, we set the epochs to 1000, the warmup epochs to 5, the scheduler to cosine, the mixup to 0.8, the smoothing to 0.1, the weight decay to 1e-4, the batch size to 64, the optimizer to SGD, the learning rate to 0.01 and the warmup learning rate to 0.0001 for all datasets. Fine-tuning is performed on 8 RTX-2080 GPUs.
F.4 NB201 and DARTS
For single-stage optimizers, the supernet was trained with four different seeds. The supernet with the best validation performance among these four was discretized to obtain the final model, which was then trained from scratch four times to obtain the results shown in the table. For two-stage methods, we again train the supernet four times and perform Random Search (RS) and Evolutionary Search (ES) on each one. The best model obtained across all four supernets for both methods was then trained from scratch with four seeds to compute the final results. For DrNAS, we follow the same procedure as suggested by the authors across all search spaces. To accommodate multiple training recipes, we have developed a configurable training pipeline. The configurations for DrNAS and SPOS are shown in Tables 24 and 25, respectively. We run our search on a single RTX-2080 for both DARTS and NB201, and train the DARTS architectures from scratch on 8 RTX-2080 GPUs.
@@ F.4 NB201 and DARTS
Appendix G Optimal architectures derived
G.1 DARTS
@@ G.3.1 CIFAR-10
mlp_ratio:$[4,4,4,4,4,4,4,4,4,3.5,4,4,3.5,4]$, num_heads:$[4,4,4,4,4,4,4,4,4,4,4,4,4,4]$, num_layers: 14 embed_dim: 216
@@
mlp_ratio:$[4,4,4,4,4,4,4,4,3.5,4,4,4,3.5,3.5]$, num_heads:$[4,4,4,4,4,4,4,4,4,4,4,4,4,4]$, depth: 14, embed_dim: 240
G.3.2 CIFAR-100
@@ G.4 MobileNetV3
G.5 Toy Spaces
G.5.1 Toy cell (our best architecture)
@@
G.6 Toy conv-macro (our best architecture)
@@
Appendix H Architecture Representation Analysis
@@
Centered Kernel Alignment (CKA) (kornblith2019similarity) is a metric based on the Hilbert-Schmidt Independence Criterion (HSIC). It is designed to model the similarity between representations in neural networks. In this section, we analyze the CKA between structurally identical layers in the inherited, fine-tuned, and retrained networks in the AutoFormer-T space. Specifically, the aim is to assess how similar the inherited and fine-tuned representations are to those of the models trained from scratch. Table 26 presents the average CKA values for a fixed subset of the CIFAR-10 and CIFAR-100 datasets. We find that the representations learned by the single-stage supernet are more similar to those of the architectures that are fine-tuned or trained from scratch. This observation underscores potential issues with using inherited weights from the supernet as a proxy for search in two-stage methods, as noted by xu2022analyzing.
@@
CIFAR-10 CIFAR-100 CIFAR-10 CIFAR-100
TangleNAS 0.4630 0.5853 0.5712 0.6527
SPOS+ES 0.4581 0.5793 0.5694 0.6309
SPOS+RS 0.4412 0.583 0.5797 0.6389
Table 26: CKA correlation between layers.
Appendix I Limitations and Future Work
Currently, TangleNAS is designed to optimize a single objective, such as a chosen performance metric. However, in practice, there may be multiple objectives of interest, including hardware efficiency, robustness, and fairness (dooley-neurips23). Additionally, it would be valuable to explore and apply our findings across various applications in computer vision and natural language processing.